### 18 Dec 2023 v3

# Empowering Autonomous Driving with Large Language Models: A Safety Perspective

Abstract
Autonomous Driving (AD) faces crucial hurdles for commercial launch, notably in the form of diminished public trust and safety concerns from long-tail unforeseen driving scenarios. This predicament is due to the limitation of deep neural networks in AD software, which struggle with interpretability and exhibit poor generalization capabilities in out-of-distribution and uncertain scenarios. To this end, this paper explores the integration of Large Language Models (LLMs) into the AD system, leveraging their robust common-sense knowledge, reasoning abilities, and human-interaction capabilities. The proposed approach deploys the LLM as an intelligent decision-maker for planning, incorporating safety verifiers for in-context safety learning to enhance overall AD performance and safety. We present two case studies to affirm the efficacy of our approach. We further discuss the potential usage of LLM for other AD software components including perception, prediction, and simulation. Despite the observed challenges in the case studies, integrating LLMs is promising and beneficial to reinforce the safety and performance of AD.
1 Introduction
The current mainstream of the autonomous vehicle (AV) software pipeline includes several crucial modules, organized in a stack manner. Initially, the perception module consumes sensor data from Lidar, radar, and camera to provide object detection outcomes (Feng et al., 2020). Subsequently, the prediction module anticipates the high-level intentions and low-level trajectories of other relevant entities based on scene analysis (Jiao et al., 2023). The planning module then makes decisions on the vehicle’s behavior, such as lane-keeping, gap tracking, and lane changing, followed by the computation of trajectory-level waypoints for precise control tracking (Liu et al., 2023b). Control inputs are applied to the ego vehicle, evolving it to transition to a new state in accordance with its physical dynamics and constraints.
Deep neural networks (DNNs) have become integral to perception and prediction within the AV pipeline, with a growing interest in their application in planning and control. However, the black-box nature of DNNs, along with their inherent uncertainties from learning algorithms, presents challenges in ensuring the safety of closed-loop AV systems. These challenges are exacerbated by the generalizability issue faced by DNNs and the prevalence of long-tail driving scenarios not covered during training (Fu et al., 2023). Because of these, there is a lack of public confidence in AV due to the uncertain safety issues and most importantly the fact that the driving behaviors made by AV are uninterpretable without reasoning to the user.
In response to these challenges, researchers and engineers in the AV industry are exploring the potential of Large Language Models (LLMs) (Touvron et al., 2023; OpenAI, 2020; Devlin et al., 2018) for their ability for human interaction, adept reasoning capabilities, and comprehensive knowledge, particularly in handling long-tail driving scenarios for autonomous driving (Yang et al., 2023). Nevertheless, the practical integration of LLMs into the AV software pipeline for safety purposes remains an open question. Therefore, this research paper delves into the application of LLMs in autonomous driving from a safety perspective, highlighting its implementation through a couple of illuminating case studies.
Figure 1 shows our proposed possible integration of LLMs for different modules in the AV software pipeline, from a safety perspective. As a safety-critical system, we equip the AV with a safety verifier for the proposed control input generated from the software stack with assistance from LLMs. The verifier returns safety-checking results to LLM for in-context safety learning which could affect the outputs from different components in various ways. In this paper, we conduct two case studies to leverage LLM as a behavior-level decision-maker which interacts with a high-level predictor for evaluating the intention and aggressiveness of other agents, and with the low-level trajectory planner and safety verifier. These case studies show that LLM can improve system performance while achieving safety assurance through in-context safety learning. We hope this paper can provide the AV community with a comprehensive safety standpoint to explore and evaluate the usage of LLM in their AV software stack.
This paper is organized as follows. We introduce related works in Section 2. Section 3 and Section 4 show our two case studies for integrating LLM as an intelligent safety-aware behavioral decision-maker with a safety verifier for contextual safety learning. Section 5 discusses the possible integration of LLM for other components including perception, prediction, and simulation in the AV system for safety purposes. Section 6 concludes the paper.
2 Related Works
2.1 LLM for Autonomous Driving
The integration of LLMs such as GPT-3 (OpenAI, 2020) into AD has garnered significant attention in recent years, revolutionizing natural language understanding and enhancing the capabilities of self-driving vehicles (Wayve, 2023). We list the related literature from different perspectives as follows.
Human-Oriented: One direct application is enabling human-vehicle interaction through natural language. LLMs have been leveraged to interpret, respond, and provide suggestions in natural language to human riders and drivers (Zhang et al., 2023; Wayve, 2023; Xu et al., 2023). These models generate user-friendly natural language narrations that assist human driving for decision-making and improve the interpretability of AD systems by explaining driving behaviors. Recent works have gone beyond interaction and employed LLMs to learn human driving behaviors and trajectory data through chain-of-thoughts (Wei et al., 2022; Jin et al., 2023b). This approach enables the LLM driver to behave like humans to solve complex driving scenarios and even allows LLMs to function directly as motion planners (Mao et al., 2023).
Perception, Prediction, and Planning (Decision-making): The reasoning, interpretation, memorization, and decision-making abilities of LLMs contribute to solving long-tail corner cases, improving generalizability, and increasing the interpretability of AD systems. Specifically, there is a growing interest in integrating LLMs into the planning (decision-making) module, which significantly improves user trust and generalizes to various driving cases (Jin et al., 2023a). This integration is achieved through fine-tuning pre-trained LLMs (Liu et al., 2023a) or by prompt engineering with chain-of-thought, which usually enable the AD motion planner to process multilabel inputs, e.g., ego-vehicle information, maps, and perception results (Wen et al., 2023; Cui et al., 2023; Fu et al., 2023; Mao et al., 2023). Leveraging the few-shot learning ability and the capability of understanding heterogeneous inputs of LLMs, researchers are exploring their potential in the perception module to enable self-aware perception, fast and efficient adaptation to changing driving environments, including tracking, detection, and prediction (Malla et al., 2023; Radford et al., 2021; Wu et al., 2023; Ding et al., 2023). Zhou et al summarizes the state of art works in this field (Zhou et al., 2023). In addition to these approaches, Keysan et al. (Keysan et al., 2023) demonstrate that combining a text encoder from a pre-trained LLM with an image encoder achieves superior performance in trajectory prediction compared to using a single encoder in isolation.
Nevertheless, the aforementioned references fail to address safety concerns associated with the utilization of the LLM in AD. In our paper, we prioritize safety first under the context of LLM, a perspective evident in our case studies. This emphasis on safety is imperative due to the safety-critical nature inherent in AD. We take a safety-focused approach, allowing LLM decisions to directly formulate safety constraints for low-level Model Predictive Control (MPC) in the face of prediction uncertainties. Our case studies align closely with the LanguageMPC approach (Sha et al., 2023). In their work, the authors employ LLMs as a decision-making element for AD scenarios. They seamlessly converted LLM decisions into the mathematical representations needed for the low-level controllers, MPC, through guided parameter matrix adaptation, achieving a significant performance leap. However, it’s crucial to acknowledge that LanguageMPC has not been extensively validated in complex driving environments. Additionally, it lacks the ability to manage uncertainty in predictions and does not include safety analysis or optimization in its methodology.
Generation and Simulation: The introduction of LLMs has also dramatically transformed the process of collecting driving data. LLMs’ generative capabilities have facilitated the acquisition of complex driving data samples, which were previously difficult to gather due to certain environmental constraints. The diffusion model, a method that has recently reached significant success in the text-to-image domain, has become increasingly popular (Sohl-Dickstein et al., 2015; Ho et al., 2020). Some efforts has been put into the area of generating the driving scenarios using diffusion models (Li et al., 2023; Gao et al., 2023; Wang et al., 2023a; Hu et al., 2023; Zhong et al., 2023). CTG++(Zhong et al., 2023) introduces a scene-level model conditioned by language instructions, generating realistic and controllable traffic simulation environments for AV development.
2.2 Safety Verification for ML-based Cyber-Physical Systems
Our work is related to the rich literature on safety verification for ML-based cyber-physical systems where AD systems are representative. Safety verification, in general, can be categorized into two groups: 1) explicit reachable set computation (Wang et al., 2023d; Huang et al., 2022; Ivanov et al., 2021; Kochdumper et al., 2023; Goubault and Putot, 2022; Schilling et al., 2022; Huang et al., 2019) and 2) inexplicit reachable set evaluation, such as barrier certificate (Prajna, 2006; Wang et al., 2023b), control barrier function (Ames et al., 2019), forward invariance (Wang et al., 2020; Chen et al., 2018), etc. There have been emerging works for integrating verification modules into the control learning or reinforcement learning for safety-assured autonomy (Dawson et al., 2022; Wang et al., 2023c, b; Zhan et al., 2023; Jin et al., 2020). Our paper follows a similar idea where we develop in-context safety learning for LLM to make safer decisions with a safety verifier in the software pipeline.
3 LLM as Behavior Planner: LLM Conditioned Adaptive MPC for Planning with Safety Assurance
As mentioned, LLM has the ability to reason and infer behavior-level decisions for autonomous driving via prompt engineering. Here we conduct a case study for LLM as a behavior planner, as shown in Figure 2. Next, we introduce the components of this case study as follows.
Environment and System: We developed a highway-driving environment by using the open-source HighwayEnv Leurent (2018). As shown in Figure 4, we consider a one-way three-lane driving scenario. We assume that the vehicle dynamics is known and available to MPC, which can be expressed as where with denote longitudinal position, lateral position, longitudinal speed, and lateral speed, respectively. The continuous control input to the ego vehicle includes acceleration and steering signal. denotes the bicycle model dynamics (Jiao et al., 2023).
Input and Output of LLM: We call OpenAI GPT-4 API as our LLM driver agent. We input the text description of the surroundings within a specific perception range including their relative position (such as "the car is driving in front of the ego on the right lane" or "the car is driving behind the ego in the middle lane"), their relative speed (such as "the car is driving faster/slower than the ego"), the estimation of time to the collision to other agents, along with other vehicle’s intention predictions. The output of the LLM decision maker is constrained to select a target lane for lower level MPC (such as "Middle Lane, Left Lane, Right Lane") with the reasoning.
Prediction Module: The prediction module on the AV predicts the future state of surrounding car at time step . To be realistic and considering uncertainties, we assume the predicted position results are intervals on a specific time step, i.e, instead of , we now have and . We assume the position intervals contain the ground truth of the surroundings in the future, i.e., , . Because of the receding horizon nature of MPC, we need to call the prediction module to get the prediction results for safety constraint formulation in MPC.
LLM Conditioned Adaptive MPC for Trajectory Planning: Before introducing the LLM conditioned MPC, we first show that a naive formulation of MPC for trajectory planning is to solve the following optimization problem
where is an indicator function that determines which lane the car is driving on by its lateral position , specifically denotes "Left", "Middle", and "Right". The objective function aims to maximize the performance (longitudinal position or speed) with minimal control inputs.
To reduce the complexity, we leverage the reasoning ability and common sense knowledge of LLM to decide which lane to drive for the MPC, by providing the scene text description to LLM and asking for a decision that relaxes the constraints in MPC. Specifically, at time step , our LLM conditioned MPC tries to solve the following optimization problem
The problem 1 is harder to solve than problem 2. The increased complexity originates from the constraint , where is undetermined and can choose from . Therefore problem 1 is a mixed integer nonlinear programming problem. In practice, this problem is often infeasible, as observed in our case studies. With the decision from LLM by its knowledge, we remove the integer decision variable in problem 2 and thus it is easier to solve. Our approach shares a similar philosophy of hierarchical MPC as introduced in (Huang et al., 2016) where we decompose a hard trajectory planning into a two-phase problem that is easier to solve.
In-context Safety Learning with Verifier: For safety purposes, control input to the ego vehicle has to go through a verification module (verifier) to conduct safety verification and provide the verification result back to the LLM to reevaluate the behavior decision. The verifier could be in the form of reachability analysis (Wang et al., 2023d), barrier theory (Wang et al., 2023c), and/or feasibility of low-level MPC with safety constraints as in problem 2. In this case study, because of the assumption of the interval-based position prediction of the surroundings and the safety constraint in the MPC, if the MPC problem 2 is feasible, we then feedback “the verifier is happy with the proposed Lane(LLM)” to LLM. Otherwise, infeasible MPC leads to potential unsafe planning waypoints which we feedback to LLM to reevaluate the scenario and regenerate another behavior. The in-context safety learning of LLM is shown in Figure 3.
Failsafe Mode: It is possible that regenerated behavior or all behaviors are still infeasible for the low-level MPC and thus safety cannot be assured. In this case, we design the AV system switch to a failsafe mode, where the behavior is to keep the current lane and brake to keep a minimal distance from the front leading car. We put safety as the highest priority and thus the ego vehicle can apply a conservative hard brake to reduce its speed for the lane keeping.
Experiments Analysis: We compare our approach with the state-of-the-art DriveLikeAHuman (Fu et al., 2023) because both approaches are testified in the HighwayEnv simulator. We add the interval-based prediction uncertainty to the DriveLikeAHuman framework, the same as our setting, for a fair comparison. We simulate 300 control steps in one test episode where DriveLikeAHuman has to call an LLM decision maker at every single step while our approach calls LLM once after every 5 control steps by MPC. The maximum velocity is set to . We conduct five trials for each method and record their performance as below. Our case study can be accessed at this Github repo.
Safety: Safety (Non-collision) is assured and observed in our simulations with 1500 total control steps and more than 300 LLM decision-making processes. This is due to the assumption that interval-based prediction encoded in the constraints of MPC covers the ground truth of future locations of the surroundings. Therefore, a feasible MPC solution and the failsafe mode when infeasible assure safety in our framework. On the contrary, with DriveLikeAHuman, we found collisions happened in 4 out of 5 trials around steps while the one left test terminated due to an LLM calling error after 77 steps. This is because 1) DriveLikeAHuman uses a simple heuristic rule for safety checking with low-level PID control and does not consider vehicle dynamics and constraints and thus cannot fully capture the safety 2) DriveLikeAHuman cannot handle the prediction uncertainty in its safety checking heuristics.
Performance: We measure the longitudinal speed average and standard deviation as performance metrics for all the simulated control steps. Our approach obtains while the baseline only gets . This is because we maximize the longitudinal location (speed) in the objective function of our LLM-conditioned MPC.
Latency: We measure the latency per control step as the third metric. The latency of our approach includes the communication latency with ChatGPT every 5 control steps and the timing of MPC every step while the baseline spends most of the time on the chain-of-thought process with ChatGPT per control step. Our approach on averages takes seconds and DriveLikeAHuman takes seconds.
Behaviors: We observed that LLM in our approach can intelligently switch between lane keeping and lane changing to maximize performance. It can also abort lane-changing when the low-level MPC is not feasible during the lane-changing behavior, as shown in Figure 4.
Problem Observed: We have observed instances where the LLM exhibits unnecessary lane changes even when both lanes are clear and safe. The rationales behind these decisions range from "the middle lane offers more operational space for maneuvering" to "the target lane provides a better balance between efficiency and safety." While GPT-4 demonstrates commendable performance, as highlighted earlier, GPT-3.5-turbo tends to grapple with distinguishing between "longer" and "shorter" time to collision (TTC), resulting in suboptimal and incorrect decisions, particularly in later iterations. Despite initially grasping TTC accurately in the early runs, GPT-3.5-turbo later contradicts itself. In summary, our observations indicate that GPT-4 consistently outperforms GPT-3.5-turbo in behavioral decision-making for autonomous driving.
4 LLM as Interactive Decision Maker: Interactive Planning by Behavior Prediction and State Machine
As with most existing works on LLM for AD, our previous case study focuses on such one-step planning or single-frame decision-making. We can further improve the performance and safety of LLM for driving tasks by explicitly considering the ego vehicle’s high-level behavior patterns (controlled by LLM) and the prediction and interaction with surrounding agents in multiple consecutive steps. Based on the first case study, the highlights of this second case study are as follows.
State Machine as Behavior Guidance for LLM Decision Maker: In practical driving situations, executing maneuvers like lane changes or unprotected left turns involves several steps. Vehicles must adhere to established rules or patterns to interact effectively with nearby traffic and accomplish these tasks. This necessitates that LLMs conform to specific behavioral constraints. We propose to integrate a state machine into the LLM through in-context learning or few-shot fine-tuning. This approach aims to ensure adherence to these essential regulations and established patterns while maintaining sufficient flexibility in decision-making.
The state-machine-conditioned LLM can enhance the safety and interpretability of the decision-making process in several aspects. First, the state machine can constrain the decision space and simplify the dependency among time-series decisions. Humans can trust the decision-making pipeline if the LLM ensures the transition between states is safe. Second, we add some intermediate/interactive states into the state machine design to help the LLM better understand other vehicles’ behavior. During these states, the LLM can proactively interact with the surrounding vehicle (e.g. the following vehicle on the target lane when changing lanes) but still ensure safety.
In Figure 5, we present our pipeline for interactive lane changing using LLM as the decision-maker. The framework is centered around the state machine which defines the basic behavior pattern of our LLM. The memory stores important past information about surrounding vehicles, helping the LLM make predictions of their intentions. The reflection module is to monitor the LLM and make sure the transition is valid from state to state and to give feedback to the LLM for in-context learning when the LLM violates transition constraints. The LLM determines transitions based on predefined rules and inferred information. The transition involves several checks:
State Check: The selected state must be valid as per a predefined state machine graph.
Safety Check: The LLM evaluates the possibility of collision if it takes certain actions transiting to the next state. In this particular example of interactive lane changing, the time-to-collision (TTC) is applied to ensure the proposed state won’t lead to a collision. The LLM will compare the TTC against a set threshold.
Prediction Check: The LLM predicts the intentions of nearby vehicles based on their historical behaviors in past multiple frames in the memory modules. If the LLM deems a surrounding vehicle too aggressive or uncooperative, it’s unsafe to proceed with the maneuver. The LLM can interact with the surrounding vehicles in different manners given their different predicted behavior patterns.
Reflection Module: State and safety checks are stringent requirements in the decision-making process. To ensure compliance with these requirements, a reflection module monitors state transitions. This module corrects and provides feedback to the LLM, facilitating in-context learning, especially when decisions breach these strict constraints. For behavior prediction, the reflection module enforces no constraints to the intention estimation - the LLM independently and flexibly assesses the intentions of surrounding vehicles, categorizing them as either aggressive or cooperative.
Intention Prediction Module: Unlike the prediction for MPC, the intention prediction is to estimate the high-level behavior patterns of the surrounding vehicle, which doesn’t need to be very detailed but is important for interaction. We define the potential intention of surrounding agents as cooperative and aggressive. We use the time-to-collision (TTC) of surrounding vehicles to measure their aggressiveness levels. At every planning step, the LLM decision-maker will extract the surrounding vehicles’ TTCs with the past 3 steps and predict their corresponding intentions. We will give several human-labeled demonstrations when setting up the LLM.
Experimental Analysis: In the experiment, we demonstrate how the proposed framework enables the LLM to conduct interactive and safe motion planning in consecutive steps based on a predefined state machine. As shown in Figure 7 and Figure 6, we notice that the LLM can ensure safety in the behavior level by adhering to the three constraints. The LLM continually estimates the surrounding vehicle’s intention and the estimation is further confirmed during the proactive interaction state - "Attempt". After LLM makes its decision, the behavior safety checker and MPC will ensure the feasibility and safety in the behavior level and the control level, respectively.
5 Discussion: LLM as Other Roles for Safety
In previous case studies, we first naively leveraged LLM as a one-step behavior decision-maker and then further developed it into interactive planning with state machine guidance and behavior prediction for surrounding agents in multiple steps. Observing improved performance and safety in the case studies, we are going to discuss the possible usage of LLMs for other components in the AD software pipeline, as shown in Figure 1. We directly ask the question for ChatGPT-3.5 (e.g., prompt as “How can a large language model assist the perception module for safer autonomous driving?”) and summarize its responses below.
5.1 LLMs in Perception
1) Multimodel Fusion: It is possible to consider multimode infusion with language input. By integrating information from both sensor data and language input, the perception module can create a more comprehensive understanding of the environment. This multimodal fusion enables the system to make more informed safer decisions by considering both visual information and contextual cues provided by natural language. 2) Semantic Object Recognition: LLMs can assist in recognizing and understanding objects in the environment based on their semantic context of safety. For instance, if a passenger says, “Watch out for the cyclist ahead”, LLMs can understand this information to prioritize and adapt the behavior accordingly, enhancing safety. 3) Adaptive Object Detection: LLMs can provide information that helps the perception module adapt its object detection algorithms based on specific scenarios. For example, if LLMs understand that the vehicle is in a construction zone, they can convey this information to the perception module, prompting the system to be more cautious and attentive to potential hazards. 4) Human-Interacted Perception: The language model can generate annotations or labels for objects detected by the perception module in the environment with natural language which could increase the user’s trust. Furthermore, LLMs can enhance the perception module’s ability to recognize and respond to human-related cues including the ones related to safety. 5) Reducing False Positives: The language model can help filter out false positives in object detection by considering contextual information. If a sensor generates a potential obstacle, the language model can verify its validity based on the overall context, reducing the likelihood of unnecessary interventions.
5.2 LLMs in Prediction
1) Natural Language Inputs for Contextual Awareness: The language model in the prediction module can process natural language inputs (possibly from perception) to understand and infer the potential intentions of other drivers. For example, if the perception model or human user interprets "heavy traffic ahead," the prediction module with LLMs can understand it and adjust its expectations and predictions accordingly for safer operation. 2) Human-Centric Predictions: Language understanding can help the prediction module make more human-centric predictions by considering factors such as hand gestures, turn signals, or spoken commands from other drivers. This allows the autonomous vehicle to anticipate and respond to human behaviors more effectively, improving AV safety. 3) Behavioral Evaluation: The language model can assist in evaluating the driving behaviors and aggressiveness of surrounding cars by processing information such as traffic reports or user-generated input. This helps the prediction module adjust its predictions based on the perceived driving styles of other vehicles. 4) Learning from Historical Data: Due to the memorization ability, the language model can facilitate learning from historical data by interpreting user feedback or incorporating information from past interactions. This adaptive learning approach allows the prediction module to improve its predictions over time based on real-world experiences.
5.3 LLMs in Simulation
1) User Specific Scenario Generation and Variation: The language model can generate natural language descriptions of diverse driving scenarios by user input for safety concerns, allowing the simulation module to create a wide range of realistic and challenging situations for testing and training in a safety perspective. This helps in ensuring that the autonomous system is well-prepared for various real-world conditions. 2) Human-Like Interaction: The language model can simulate human-like interactions by generating realistic communication between simulated drivers, pedestrians, and other entities. This enhances the realism of the simulation, allowing the autonomous system to practice responding to natural language cues and gestures for safety purposes. 3) Simulation Annotation and Analysis: The language model can assist in annotating simulation data by generating descriptions or labels for different events and entities, which further the AV development. 4) Fault and Unsafe Edge Case Analysis: The language model can aid in summarizing and generating natural language descriptions of simulated faults and unsafe edge cases in the simulation environment. This helps the developers to better understand and improve the AD systems.
6 Conclusion
In conclusion, our presented framework explores the integration of an LLM in an autonomous driving system, fortified by a safety verifier within the software pipeline to facilitate in-context safety learning. The versatility of LLMs is leveraged across perception, prediction, and planning, thereby augmenting overall system safety, as extensively elaborated in this paper. Through two compelling case studies, we demonstrate the efficacy of employing LLMs as behavior decision-makers, showcasing notable enhancements in both performance and safety. We intend to broaden the safety perspective within the research community concerning the utilization of LLMs in the domain of autonomous driving. We believe that this paper catalyzes further exploration and innovation in ensuring the safety, reliability, and interpretability of autonomous vehicles.